Please help with these questions

You must show all calculation steps to receive a full credit. Number of questions depends on chapte given for each exam-so it varies. 1. Which one of the following is not an electromagnetic wave? A) ultraviolet B) infrared C) radio waves D) sound waves E) gamma rays 2. 2) Which one of the following lists gives the correct order of the electromagnetic spectrum from low to high frequencies? A) radio waves, infrared, microwaves, ultraviolet, visible, x-rays, gamma rays B) radio waves, ultraviolet, x-rays, microwaves, infrared, visible, gamma rays C) radio waves, microwaves, infrared, visible, ultraviolet, x-rays, gamma rays D) radio waves, microwaves, visible, x-rays, infrared, ultraviolet, gamma rays E) radio waves, infrared, x rays, microwaves, ultraviolet, visible, gamma rays 3. 8) For an electromagnetic wave in free space having an electric field of amplitude Eand a magnetic field of amplitude B, the ratio of B/E is equal to A) c B) c2 c) 1/c D) 1/c2 E) 4. What is the wavelength used by a radio station that broadcasts at a frequency of 920 kHz? (c -3.00 x 108 m/s) A) 22.6 m B) 226 m C) 326 m D) 175 m E) 276 m S. The frequency of a microwave signal is 9.76 GHz. What is its wavelength? (c 3.00x 108 m/s) A) 3.07 cm B) 2.07 cm C) 1.07 cm D) 5.07 cm E) 4.07 cm
Answer #1

1. D)

Electromagnetic waves are generated by periodic and simultaneous oscillations of the electric and magnetic field component and are characterised as transverse waves. Sound waves are not electromagnetic waves and are characterised as longitudinal waves.

2. C)

Out of the given options only the option C) arranges the electromagnetic waves in increasing order of frequencies and in decreasing order of their wavelength.

3. C)

E is the electric field vector and B is the magnetic field vector. Their ratio gives the speed of light (c). Mathematically we may write,

E/B = c or B/E = 1/c

4. C)

Relation between frequency (f), wavelength (lambda) and speed of electromagnetic waves is given as

or X=c/f

In the given problem,

c = 3 imes 10^8 m/s f = 920 KHz = 920 imes 10^3 Hz

using the values in the equation for lambda,

lambda = rac{3 imes 10^8}{920 imes 10^3} = 3.26 imes 10^2 = 326 m

5. A)

In the problem,

c= 3 × 108 m/s f = 9.76 GH: 9.76 × 109 Hz

using the same relation for lambda to get,

lambda = c/f = rac{3 imes 10^8}{9.76 imes 10^9} = 0.3 imes 10^{-1} m = 0.0307 m = 3.07 cm (since 1 m = 100 cm)